# Unified Source-Free Domain Adaptation

### v1

@@ 3.4.1 External Causality Discovery
To obtain external causality for a target task, we explore the potential of recent large multimodal ViL models such as CLIP [11]. This is because they have experienced a vast amount of knowledge in pretraining, and exhibited general ability in many different perception tasks.
Specifically, we accomplish this idea by encoding $S_{e}$ into the prompt context $\boldsymbol{v}_{e}$ of a ViL model $g_{\theta}$ efficiently. That is, our learning target is $\boldsymbol{v}_{e}$, which is the explicit expression of $S_{e}$. We start by formulating the $S_{e}$ discovery as maximizing the correlation between a random variable representing $S_{e}$ and the prediction of the target model. Then, the original formulation is converted to a self-supervised information bottleneck problem with theoretical guarantee, followed by the deep learning instantiation.
With the target samples $\mathcal{X}$ and learnable $\boldsymbol{v}_{e}$, we obtain the prediction of $g_{\theta}$ as $\hat{\mathcal{P}}=\{\hat{\boldsymbol{p}}_{i}\}_{i=1}^{n}$. Meanwhile, we obtain the logits of target model $f_{\theta_{t}}$ for the same data as ${\mathcal{A}}=\{{\boldsymbol{a}}_{i}\}_{i=1}^{n}$. We introduce three random variables, $V$, $Y$, and $Z$, following the probability distributions of $\boldsymbol{v}_{e}$, $\hat{\mathcal{P}}$ and ${\mathcal{A}}$, respectively. In LCFD, we further write $Z=\{Z_{e},Z_{i}\}$ where the latent random variables $Z_{e}$ and $Z_{i}$ represent the external $S_{e}$ and the internal $S_{i}$ causal factors, respectively.
In information-theoretic view, for causality relationships we have to ensure the correlation between ${Z}_{e}$ and $V$. This could be considered an approximation with computational ease. Formally, discovering $S_{e}$ needs maximizing the following mutual information:
@@
where function $I(\cdot,\cdot)$ computes mutual information of two random variables.
Consider that prompt $V$ is specific to the ViL model $g_{\theta}$, we further transform the above formula as
@@
where $\phi(\cdot)$ is softmax function; so that the two variables under correlation maximization reside in similarly semantic spaces. Theoretically, we prove that this transformed is a low bound of the original, thus valid in terms of optimization.
First, we have the following Lemma 6 (see the proof in Appendix).
@@
The ViL model $g_{\theta}$ is indeed compressing, as it maps a high-dimensional image to a low-dimensional category vector. Letting $Z_{1}=Z_{e}$, $X_{1}=X$, $Y_{1}=Y$ and $f_{1}=\phi\left(g_{\theta}(\cdot)\right)$, we have $I({{Z_{e}},V})\geq I({{Z_{e}},Y})$ according to Lemma 6.
Figure 4: Realising our self-supervised information bottleneck w.r.t. Theorem 8. The item $\rm{T}_{1}$ is estimated by variational information maximization with Gaussian distribution assumption $\mathcal{N}(Z,\Sigma_{Z})$. The item $\rm{T}_{2}$ is computed by mutual information definition where the function $f_{w}$ integrates the min-max optimization, using covariance matrix $\Sigma_{Z}$ from $\rm{T}_{1}$ as the weighting parameters $w$.
Self-supervised information bottleneck. In Eq. (5), $Z_{e}$ is hidden part of the observable logtis along with non-causal and internal causal factors. To facilitate this learning process of $Z_{e}$, we derive a self-supervised information bottleneck algorithm without prior assumption on the distribution [37]. In general, information bottleneck conducts min-max optimization to squeeze out the essentials w.r.t. a task. We introduce a random variable $Z{{}^{\prime}}$ (i.e., the bottleneck) for the ViL’s logits $\widehat{\mathcal{O}}=\{\hat{\boldsymbol{o}}_{i}\}_{i=1}^{n}$ between $V$ and $Y$, thereby forming a bottleneck style expression as:
@@
where $Z$ is the logit variable of target model $f_{\theta_{t}}$ which is observed. Note, we do not include the item $I\left({Z{{}^{\prime}},Y}\right)$ in minimization due to lacking of supervision signal like ground-truth labels as used in conventional information bottleneck [38], i.e., a self-supervised formulation. However, we theoretically prove that Eq. (7) is an upper bound as below thus optimizationally valid.
@@
Suppose that there are five random variables $Z$, $V$, $Z{{}^{\prime}}$, $Y$ and $Y{{}^{\prime}}$. Among them, $Z$ represents the target domain knowledge; $V$, $Z{{}^{\prime}}$ and $Y$ express the input instances, an intermediate features of the ViL model and predictions, respectively; $Y{{}^{\prime}}$ depicts a pseudo-label that has a functional relationship $f_{w}$ ($w$ is learnable parameters) with $Z{{}^{\prime}}$. When $f_{w}$ is reversible and uncompressed, we have this upper bound
@@
The proof is given in Appendix LABEL:sec-thm. Theorem 8 suggests that the desired latent factors refinement is conditionally equivalent to a self-supervised information bottleneck, when pseudo-label $Y{{}^{\prime}}$ is generated from the intermediate ViL feature $Z{{}^{\prime}}$ without information loss.
We realize the proposed information bottleneck using variational information maximization [39]. The key is to construct the function $f_{w}$ with three properties: (i) The output is a probability distribution, e.g., including a softmax operation $\phi$. (ii) No information loss, e.g., no zero weights in $w$ for weighting based functions. (iii) The weighting parameters $w$ are constrained to be relevant to $I(Z{{}^{\prime}},Z)$. The instantiation of $T_{1}$ and $T_{2}$ are elaborated below.
@@
where $q({Z{{}^{\prime}}}|Z)$ is variational distribution approximating the real distribution $p({Z{{}^{\prime}}}|Z)$, ${\rm{KL}}(\cdot,\cdot)$ is ${\rm{KL}}$-divergence function, and $H({Z{{}^{\prime}}})$ is a constant. In practice, we model $q({Z{{}^{\prime}}}|Z)$ as a Gaussian distribution $\mathcal{N}(Z,\Sigma_{Z})$ with mean $Z$ and diagonal covariance matrix $\Sigma_{Z}$. Then, the loss for minimizing $T_{1}$ can be formulated (see $\rm{\mathbf{T}}_{1}$ in Fig. 4):
@@
Considering that the ViL’s prediction $Y$ and the pseudo-label $Y{{}^{\prime}}$ are probability distributions, we compute the second term $\rm{T_{2}}$ by the original mutual information [41]: $I(Z,Z{{}^{\prime}})={p(Z,Z{{}^{\prime}})}\mathord{\left/{\vphantom{{}{}}}\right.\kern-1.2pt}{[p(Z)p(Z{{}^{\prime}})]}$. The corresponding loss function is
$\begin{split}&\mathcal{L}_{\rm{PMI}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right)=\frac{1}{n}\sum_{i=1}^{n}{I}\left(\hat{\boldsymbol{p}}_{i},{\boldsymbol{p}}^{\prime}_{i}\right),\\ &\hat{\boldsymbol{p}}_{i}=\phi(\hat{\boldsymbol{o}}_{i}),~{}~{}{\boldsymbol{p}}^{\prime}_{i}=f_{w}(\hat{\boldsymbol{o}}_{i})\end{split}$ (11)
where $f_{w}$ is a weighting-based function we construct as follows.
@@
$\begin{split}{\boldsymbol{p}}^{\prime}_{i}=f_{\left[w=\Sigma_{Z}^{i}\right]}\left(\hat{\boldsymbol{o}}_{i}\right)=\phi\left(\frac{1}{\Sigma_{Z}^{i}}\right)\odot\phi\left(\hat{\boldsymbol{o}}_{i}\right).\end{split}$ (12)
Combining Eq. (10), (11) and (12) together, our self-supervised objective for external causality discovery is formed as
$\begin{split}&\min_{{\boldsymbol{v}_{e}},\Sigma_{Z}}\mathcal{{L}}_{\rm{EC}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right)\\ &=\mathcal{{L}}_{\rm{PMI}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right)-\alpha\mathcal{{L}}_{\rm{VMI}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right).\end{split}$ (13)
Remark. Our design differs from existing information bottleneck in two aspects: While previous information extraction [42] is confined to a single model, ours is of cross-model. Our formulation is self-supervised, in contrast to previous supervised designs.
Algorithm 1 Training LCFD
@@
1:  Initialize the target model $f_{\theta_{t}}$ by $f_{\theta_{s}}$
2:  Initialize $\boldsymbol{v}_{e}$ with a fixed template
3:  for iter $k=1$ to ${M}$ do
4:     Sample a mini-batch $\mathcal{X}_{b}$ from $\mathcal{X}$.
5:     Discover external causal factors $S_{e}$: update $(\boldsymbol{v}_{e},\Sigma_{Z})$ by optimizing $\mathcal{{L}}_{\rm{EC}}\left(\mathcal{X}_{b};\boldsymbol{v}_{e},\Sigma_{Z}\right)$, fixing $\mathcal{{L}}_{\rm{IC}}$.
6:     Convert updated $\boldsymbol{v}_{e}$ to target pseudo-labels for this batch $\mathcal{X}_{b}$, i.e., obtain ${\mathcal{Q}}_{b}=\phi(g_{\theta}(\mathcal{X}_{b},\boldsymbol{v}_{e})$.
7:     Discover internal causal factors $S_{i}$: update $f_{\theta_{t}}$ by optimizing $\mathcal{{L}}_{\rm{IC}}$, fixing $\mathcal{{L}}_{\rm{EC}}$.
8:  end for
9:  return: The adapted model $f_{\theta_{t}}$.
3.4.2 Internal Causality Discovery
In phase-2, we subsequently discover the internal causal factors $S_{i}$ by updating the target model. That is, we encode $S_{i}$ into the target model as it is latent in nature. This is achieved using the pseudo-labels predicted by the ViL model and prompt context in a self-supervised learning manner.
Formally, denote ${\mathcal{Q}}=\phi(g_{\theta}(\mathcal{X},\boldsymbol{v}_{e}))$ the pseudo-labels of the target data $\mathcal{X}$ generated by the ViL model $g_{\theta}$, using $\boldsymbol{v}_{e}$ updated in phase-1, and ${\mathcal{P}}=\phi(f_{\theta_{t}}(\mathcal{X}))$ the predictions of $\mathcal{X}$ by the target model $f_{\theta_{t}}$. We design the objective function for discovering $S_{i}$ as
$\footnotesize{{\cal L}_{{\rm{IC}}}}\mathop{=}\min_{{\theta_{t}}}\underbrace{H({\cal P})+\tau\sum\limits_{c=1}^{C}{{{{\rm{KL}}}}}\left({\left.{{\varrho_{c}}}\right\|{1\over C}}\right)}_{{{\cal L}_{{\rm{UN}}}}}+~{}\sigma\underbrace{{{\cal L}_{{\rm{SCE}}}}\left({{\cal P},{\cal Q};{\theta_{t}}}\right)}_{{{\cal L}_{{\mathop{\rm PL}\nolimits}}}}$ (14)
where ${H}({\mathcal{P}})\!=\!-\sum_{i=1}^{n}\boldsymbol{p}_{i}\log\boldsymbol{p}_{i}$ estimates the information entropy, $\frac{1}{C}$ stands for the uniform distribution over $C$ categories, ${\varrho}_{c}=\frac{1}{n}\sum_{i=1}^{n}{p}_{i,c}$ is empirical label distribution, ${p}_{i,c}$ is the probability of $\boldsymbol{x}_{i}$ in the $c$-th category, ${{\cal L}_{{\rm{SCE}}}}$ is the softmax-cross-entropy loss, and $\tau$, $\sigma$ are trade-off parameters. The first term ${\cal L}_{{\rm{UN}}}$ is a regularization for encouraging prediction confidence and avoiding solution collapse [43]. Aligning with the pseudo labels estimated using the prompt context, the second item ${\cal L}_{{\rm{PL}}}$ essentially integrates the external causal factors $S_{e}$ while learning $S_{i}$.

### v4

@@ 3.4.1 External causality discovery
To obtain external causality for a target task, we explore the potential of recent large multimodal ViL models such as CLIP [42]. This is because they have accumulated a vast amount of knowledge and exhibited strong generalization abilities across a wide range of perception tasks.
Specifically, we accomplish this idea by encoding $S_{e}$ into the prompt context $\boldsymbol{v}_{e}$ of a ViL model $g_{\theta}$ efficiently. That is, our learning target is $\boldsymbol{v}_{e}$, which is the explicit expression of $S_{e}$. We start by formulating the $S_{e}$ discovery as maximizing the correlation between a random variable representing $S_{e}$ and the prediction of the target model. Then, the original formulation is converted to a self-supervised information bottleneck problem with a theoretical guarantee, followed by the deep learning instantiation.
With the target samples $\mathcal{X}$ and learnable $\boldsymbol{v}_{e}$, we obtain the prediction of $g_{\theta}$ as $\hat{\mathcal{P}}=\{\hat{\boldsymbol{p}}_{i}\}_{i=1}^{n}$. Meanwhile, we obtain the logits of target model $f_{\theta_{t}}$ for the same data as ${\mathcal{A}}=\{{\boldsymbol{a}}_{i}\}_{i=1}^{n}$. We introduce three random variables, $V$, $Y$, and $Z$, following the probability distributions of $\boldsymbol{v}_{e}$, $\hat{\mathcal{P}}$ and $\mathcal{A}$}, respectively. In CausalDA, we further write
$Z=\{Z_{e},Z_{i}\}$ (4)
where the latent random variables $Z_{e}$ and $Z_{i}$ represent the external $S_{e}$ and the internal $S_{i}$ causal factors, respectively.
In an information-theoretic view, for causality relationships, we have to ensure the correlation between ${Z}_{e}$ and $V$. This could be considered an approximation with computational ease. Formally, discovering $S_{e}$ needs maximizing the following mutual information:
@@
where $I(\cdot,\cdot)$ is mutual information of two random variables. Note, mutual information is not a prerequisite for causality discovery; instead, it is an effective design choice that enhances the discovery of external causality.
To achieve Eq. (5), we employ prompt learning [70] to accomplish the optimization. This is because it allows us to leverage CLIP without tuning model parameters. This avoids potential issues such as catastrophic forgetting and unintended distortion of the pretrained knowledge within the ViL model. Moreover, by preserving CLIP’s pretrained knowledge, prompt learning facilitates the discovery of external causality, enabling more effective integration of external knowledge into the causal learning process.
Consider that prompt $V$ is specific to the ViL model $g_{\theta}$, we further transform the above formula of Eq. (5) as
@@
where $\phi(\cdot)$ is softmax function; so that the two variables under correlation maximization reside in similar semantic spaces. Theoretically, we prove that this transformation can constitute a lower bound on the original, ensuring its validity in optimization.
First, we have the following Lemma 7 (see the proof in Appendix-A).
@@
The ViL model $g_{\theta}$ is indeed compressing, as it maps a high-dimensional image to a low-dimensional category vector. Let $Z_{1}=Z_{e}$, $X_{1}=X$, $Y_{1}=Y$ and $f_{1}=\phi\left(g_{\theta}(\cdot)\right)$, we have $I({{Z_{e}},V})\geq I({{Z_{e}},Y})$ according to Lemma 7.
Self-supervised information bottleneck. In Eq. (6), $Z_{e}$ is the hidden part of the observable logits along with non-causal and internal causal factors. To facilitate this learning process of $Z_{e}$, we derive a self-supervised information bottleneck algorithm without prior assumption on the distribution [22]. In general, information bottleneck conducts min-max optimization to squeeze out the essentials w.r.t. a task. We introduce a random variable $Z{{}^{\prime}}$ (i.e., the bottleneck) for the ViL’s logits $\widehat{\mathcal{O}}=\{\hat{\boldsymbol{o}}_{i}\}_{i=1}^{n}$ between $V$ and $Y$, thereby forming a bottleneck style expression as:
@@
where $Z$ is the logit variable of target model $f_{\theta_{t}}$ which is observed. Note, we do not include the item $I\left({Z{{}^{\prime}},Y}\right)$ in minimization due to lacking supervision signal like ground-truth labels as used in conventional information bottleneck [56], i.e., a self-supervised formulation. However, we theoretically prove that Eq. (8) admits an upper bound, thus optimizationally valid.
@@
Suppose that there are five random variables $Z$, $V$, $Z{{}^{\prime}}$, $Y$ and $Y{{}^{\prime}}$. Among them, $Z$ represents the target domain knowledge; $V$, $Z{{}^{\prime}}$ and $Y$ express the input instances, an intermediate features of the ViL model and predictions, respectively; $Y{{}^{\prime}}$ depicts a pseudo-label that has a functional relationship $f_{w}$ ($w$ is learnable parameters) with $Z{{}^{\prime}}$. When $f_{w}$ is reversible and uncompressed, we have this upper bound:
@@
The proof is given in Appendix-B. Theorem 9 suggests that the desired latent factors refinement is conditionally equivalent to a self-supervised information bottleneck, when pseudo-label $Y{{}^{\prime}}$ is generated from the intermediate ViL feature $Z{{}^{\prime}}$ without information loss.
We realize the proposed information bottleneck using variational information maximization [19]. The key is to construct the function $f_{w}$ with three properties: (i) The output is a probability distribution, e.g., including a softmax operation $\phi$. (ii) No information loss, e.g., no zero weights in $w$ for weighting-based functions. (iii) The weighting parameters $w$ are constrained to be relevant to $I(Z{{}^{\prime}},Z)$. The instantiations of $T_{1}$ and $T_{2}$ are elaborated below.
@@
where $q({Z{{}^{\prime}}}|Z)$ is variational distribution approximating the real distribution $p({Z{{}^{\prime}}}|Z)$, ${\rm{KL}}(\cdot,\cdot)$ is $\rm{KL}$}-divergence function, and $H({Z{{}^{\prime}}})$ is a constant. In practice, we model $q({Z{{}^{\prime}}}|Z)$ as a Gaussian distribution $\mathcal{N}(Z,\Sigma_{Z})$ with mean $Z$ and diagonal covariance matrix $\Sigma_{Z}$. For minimizing $T_{1}$, we adopt the Variational Mutual Information (VMI) loss (see $\rm{\mathbf{T}}_{1}$ in Fig. 4):
@@
Consider that the ViL’s prediction $Y$ and the pseudo-label $Y{{}^{\prime}}$ are probability distributions, we compute the $\rm{T_{2}}$ term by vanilla mutual information [21]: $I(Z,Z{{}^{\prime}})={p(Z,Z{{}^{\prime}})}\mathord{\left/{\vphantom{{}{}}}\right.\kern-1.2pt}{[p(Z)p(Z{{}^{\prime}})]}$. The corresponding Probabilistic Mutual Information (PMI) loss function is formed as:
$\small\begin{split}&\mathcal{L}_{\rm{PMI}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right)=\frac{1}{n}\sum_{i=1}^{n}{I}\left(\hat{\boldsymbol{p}}_{i},{\boldsymbol{p}}^{\prime}_{i}\right),\\ &\text{with}\,\hat{\boldsymbol{p}}_{i}=\phi(\hat{\boldsymbol{o}}_{i}),~~{\boldsymbol{p}}^{\prime}_{i}=f_{w}(\hat{\boldsymbol{o}}_{i}),\end{split}$ (12)
where $\sum_{c=1}^{C}{\boldsymbol{p}}^{\prime}_{i,c}=1$; $f_{w}(\cdot)$ is a weighting-based function constructed as follows.
@@
$\small\begin{split}{\boldsymbol{p}}^{\prime}_{i}=f_{\left[w=\Sigma_{Z}^{i}\right]}\left(\hat{\boldsymbol{o}}_{i}\right)=\phi\left(\phi\left(\frac{1}{\Sigma_{Z}^{i}}\right)\odot\phi\left(\hat{\boldsymbol{o}}_{i}\right)\right),\end{split}$ (13)
where $\phi(\cdot)$ is softmax function. Combining Eq. (11), (12), and (13) together, our self-supervised objective for external causality discovery is formed as
$\small\begin{split}&\min_{{\boldsymbol{v}_{e}},\Sigma_{Z}}\mathcal{{L}}_{\rm{EC}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right)\\ &=\mathcal{{L}}_{\rm{PMI}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right)-\alpha\mathcal{{L}}_{\rm{VMI}}\left(\mathcal{X};\boldsymbol{v}_{e},\Sigma_{Z}\right).\end{split}$ (14)
Remark. Our design differs from the existing information bottleneck in two aspects: While previous information extraction [27] is confined to a single model, ours is a cross-model design. Our formulation is self-supervised, in contrast to previous supervised designs.
3.4.2 Internal causality discovery
In phase-2, we subsequently discover the internal causal factors $S_{i}$ by updating the target model. That is, we encode $S_{i}$ in the target model, as it is latent in nature. Similar to capturing the external factors $S_{e}$ (Eq. (5)), this encoding can be formulated by maximizing the following conditional mutual information:
$\max_{\theta_{t}}I\left({{Z_{i}},Z_{\theta}}\right),~~{\rm{s.t.}}~\max_{\theta_{t}}I\left({{Z_{e}^{*}},Z_{\theta}}\right)$ (15)
where random variable $Z_{\theta}$ represents target model parameters $\theta_{t}$, whilst $Z_{i}$ denotes random variable of internal causal factors $S_{i}$ (Eq. (4)); $Z_{e}^{*}$ represents the currently discovered external causal factor $S_{e}^{*}$. The primary term, $\max_{\theta_{t}}I\left({{Z_{i}},Z_{\theta}}\right)$, aims to capture the internal causality $S_{i}$. The condition, $\max_{\theta_{t}}I\left({{Z_{e}^{*}},Z_{\theta}}\right)$, ensures the target model simultaneously accommodates the discovered external causality $S_{e}^{*}$.
Given ultra-high dimension and complexity of $\theta_{t}$}, combined with the hidden nature of $Z_{i}$, directly optimizing or computing Eq. (15) is intractable. To address this, we interpret the interaction between $Z_{\theta}$ and $Z_{i}$ as the inherent modeling process that transforms an input sample from pixel space into a predictive probability space. This perspective allows us to reformulate the problem as maximizing the mutual information between the target model’s input and output, both of which are observed and operational. Crucially, theoretical proofs demonstrate that mutual information between inputs and outputs is maximized when outputs are both highly confident and exhibit balanced class distributions [24]. Inspired by this theoretical finding, we maximize this originally non-computable mutual information by minimizing the following objective:
$\small\begin{split}{\cal L}_{\rm{UN}}({\mathcal{P}};\theta_{t})&=-\sum_{i=1}^{n}\boldsymbol{p}_{i}\log\boldsymbol{p}_{i}+\tau\sum\limits_{c=1}^{C}{{{{\rm{KL}}}}}\left({\left.{{\varrho_{c}}}\right\|{1\over C}}\right),\\ \text{with }&\boldsymbol{p}_{i}\in{\mathcal{P}}=\phi(f_{\theta_{t}}(\mathcal{X})).\end{split}$ (16)
In this equation, $\mathcal{P}$} represents the predictions of $n$ training samples $\mathcal{X}$ generated by the target model $f_{\theta_{t}}$. The first term minimizes the information entropy of each prediction $\boldsymbol{p}_{i}=[p_{i,c}]_{c=1}^{C}$. The KL divergence term, on the other hand, regulates the predictive balance across all $C$ categories via pulling the empirical distribution ${\varrho}_{c}=\frac{1}{n}\sum_{i=1}^{n}{p}_{i,c}$ towards a uniform distribution. The trade-off parameter $\tau$ balances these two terms.
Similarly, we convert the external causality $S_{e}^{*}$ from its prompt format $\boldsymbol{v}_{e}$ to pseudo-labels by evaluating the ViL model $g_{\theta}$: ${\mathcal{Q}}=\phi(g_{\theta}(\mathcal{X},\boldsymbol{v}_{e}))$. This conversion simplifies model optimization, as it enables the use of the conventional softmax cross-entropy loss, ${{\cal L}_{{\rm{SCE}}}}\left({{\cal P},{\cal Q};{\theta_{t}}}\right)$.
Combining these elements, we optimize Eq. (15) using the following objective function:
${\cal L}_{\rm{IC}}=\min_{\theta_{t}}{\cal L}_{\rm{UN}}({\cal P};\theta_{t})+\sigma{{\cal L}_{{\rm{SCE}}}}\left({{\cal P},{\cal Q};{\theta_{t}}}\right)$ (17)
where $\sigma$ serves as another trade-off parameter.
Algorithm 1 Training CausalDA
@@ 3.4.2 Internal causality discovery
1: Initialize target model $f_{\theta_{t}}$(by $f_{\theta_{s}}$), $\boldsymbol{v}_{e}$(by fixed template)
2:for iter $k=1$ to M{M} do
3:  Sample a mini-batch $\mathcal{X}_{b}$ from $\mathcal{X}$.
4:  Discover external causal factors $S_{e}$: update $(\boldsymbol{v}_{e},\Sigma_{Z})$ by optimizing $\mathcal{{L}}_{\rm{EC}}\left(\mathcal{X}_{b};\boldsymbol{v}_{e},\Sigma_{Z}\right)$, fixing $\mathcal{{L}}_{\rm{IC}}$.
5:  Convert updated $\boldsymbol{v}_{e}$ to target pseudo-labels for this batch $\mathcal{X}_{b}$, i.e., obtain ${\mathcal{Q}}_{b}=\phi(g_{\theta}(\mathcal{X}_{b},\boldsymbol{v}_{e})$.
6:  Discover internal causal factors $S_{i}$: update $f_{\theta_{t}}$ by optimizing $\mathcal{{L}}_{\rm{IC}}$, fixing $\mathcal{{L}}_{\rm{EC}}$.
7:end for
8:return: The adapted model $f_{\theta_{t}}$.